Vivian and Victor have the joint income of $120,000 in 2018. Three years ago they purchased 200 shares of stock at $30 a share. In 2018, they sold the 200 shares for $42 a share. What is the amount of federal income tax they owe as a result of this sale?
Answer #1

The rate at which the taxable person will have to pay on the gain of the long term capital gain is as follows;

Long-Term Capital Gains Tax Rate

Single Filers (taxable income)

Married Filing Jointly

Heads of Household

Married Filing Separately

0%

$0-$39,375

$0-$78,750

$0-$52,750

$0-$39,375

15%

$39,376-$434,550

$78,751-$488,850

$0-$461,700

$39,376-$244,425

20%

Over $434,550

Over $488,850

Over $461,700

Over $244,425

Assuming that they are married filing jointly and since their joint income is $120000 they will have to pay 15% on the gain on the capitals

Calculation of capital gain = Full value of consideration - cost of acquisition

= (200*42) -(200*30) = 8400 - 6000 = $2400

Total federal income tax on the gain = 2400 * 15% = $360